OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the curve at the given point: x^2 + y^2 = (2x^2 + 2y^2 - x)^2

OpenStudy (anonymous):

... @ (0,1/2)

OpenStudy (anonymous):

not that i tmatters anyway i'm just a little confused regarding the differentiation here

OpenStudy (anonymous):

alright give me a second to remember lol

OpenStudy (anonymous):

alright i'll upload a picture on what to do

OpenStudy (anonymous):

and an example give me a sec

OpenStudy (anonymous):

cool thanks

OpenStudy (anonymous):

like i got as far as [4x+4y*y']/[4x+4y*y'-1] = 2x^2 + 2y^2 - x

OpenStudy (anonymous):

actualy i can't... i'll just type it out what my book says GUIDELINES FOR IMPLICIT DIFFERENTIATION 1. Differentiat both sides of the equatio with respects to x 2. collect al terms involve dy/dx on the left side of the equation and move all others to right 3.factor dy/dx out of the left side of the equation 4.solve for dy/dx by dividing both sides

OpenStudy (anonymous):

by factor

OpenStudy (anonymous):

right righ tok so and

OpenStudy (anonymous):

so what do i do from what i have am i good so far

OpenStudy (anonymous):

let me get a piece of paper lol

OpenStudy (anonymous):

ama differentiate for u

OpenStudy (anonymous):

thanks k cool coco

OpenStudy (anonymous):

love it

OpenStudy (anonymous):

so is coco checking if your differentiation is right?

OpenStudy (anonymous):

one question. just verifying... for the secon part, everyting is raised to teh 2nd power?

OpenStudy (anonymous):

i'm guessing

OpenStudy (anonymous):

oh, ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ouch

OpenStudy (anonymous):

umm. i suck @ algebra... and i think i am going to mess up if i do this. but, i am still going to go ahead and square all that. can someone then tell me if i was right?

OpenStudy (anonymous):

yeah i got completely something else than he did

OpenStudy (anonymous):

the second is a chain and the first is just regular

OpenStudy (anonymous):

i got 2x+2yy'=2(4x+4yy'-1)

OpenStudy (anonymous):

=2x+2yy'= 8x+8yy'-2

OpenStudy (anonymous):

if you do implicit differentiation, then don't you do (for example, the first part): 2xdx+2ydy... then get all teh dx's and dy's on one side and divide so that dy/dx

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[2yy'-8yy'=6x-2\]

OpenStudy (anonymous):

ahh... i get it.. am so stupid! i was wondering where you got the chain rule from

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

then you can pull out y'

OpenStudy (anonymous):

i think it's\[\frac{-6x-2}{6y}\]

OpenStudy (anonymous):

the - is outside parenthesis

OpenStudy (anonymous):

\[-(6x-2)/6y\]

OpenStudy (anonymous):

it's been a while though lol

OpenStudy (anonymous):

blah wait i messed up lol

OpenStudy (anonymous):

let me do this all over agani

OpenStudy (anonymous):

i forgot something

OpenStudy (anonymous):

it doesn't help that i'm also trying to help two people

OpenStudy (anonymous):

ok, sorry guys but i am lost lol. but i wanna know how it's done. so outkast, once you did the chain rule, did you multiply or what?

OpenStudy (anonymous):

my chain rule was messed up lol it should be \[2x+2yy'=2(2x^2+2y^2-x)(4x+4yy'-1)\]

OpenStudy (anonymous):

yes that's what i got

OpenStudy (anonymous):

then what did you do

OpenStudy (anonymous):

i'd say to combine the right side

OpenStudy (anonymous):

it's going to be ugly hahah

OpenStudy (anonymous):

dive 2xdx + 2ydy with 4xdx+4ydy...?

OpenStudy (anonymous):

divide*

OpenStudy (anonymous):

i think you'd have to combine the right side which is a pain

OpenStudy (anonymous):

Hey, do you know is there anyway i can like mark this question so that i can come back and look @ it 2morrow?

OpenStudy (anonymous):

no seriously?! ouch... calyne, your proff/teacher must be coldhearted to give you such a question.

OpenStudy (anonymous):

um i think it should still be in the top left forner under notifications

OpenStudy (anonymous):

it's not that bad tho

OpenStudy (anonymous):

Oh, ok. thanks!

OpenStudy (anonymous):

\[2x+2yy'=16x^3+16y^3y'+16x^2yy'+16y^2x-12x^2-4y^2-8xyy'+2x\] \[2yy'=16x^3+16^3y'+16x^2yy'+16y^2x-12x^2-4y^2-8xyy'\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

did you get what me and coco got about my last when you used the chain rule

OpenStudy (anonymous):

what

OpenStudy (anonymous):

you use the chain rule yeah du^2/du * du/dx first thing overall ik

OpenStudy (anonymous):

what you multiply that whole thing by dy/dx too

OpenStudy (anonymous):

so whta did you get after you take dy/dx of both sides

OpenStudy (anonymous):

what did you get damn

OpenStudy (anonymous):

using the chain rule i got 2x+2yy'=2(2x^2+2y^2-x)(4x+4yy'-1)

OpenStudy (anonymous):

sorry i got 2x+2yy' = 2(2x^2+2y^2-x) (4x + 4yy' - 1)

OpenStudy (anonymous):

yeahhhh

OpenStudy (anonymous):

alright so if you combine the terms on the right

OpenStudy (anonymous):

is what i did ... what did you do after

OpenStudy (anonymous):

idk that's gay

OpenStudy (anonymous):

i tried a bunch of different stuff what was your final answer i'll tell you if it's correct nah flutter that the answer is x+1/2. is that what you got. if not we're wasting time.

OpenStudy (anonymous):

not done yet haha

OpenStudy (anonymous):

yeah i have no idea.. even wolfram alpha has omething like my end answer

OpenStudy (anonymous):

unless your teacher is crazy and solved for like y lol

OpenStudy (anonymous):

wait the answer to the differentiation is that or the tangent line

OpenStudy (anonymous):

because if it's the tangent line my implicit differentiation may be correct